OpenStudy (anonymous):

The angular Displacement x (rad) of a wheel undergoing Rotation is given by x=4t-2t^2. The Speed of particle on the rim 2s after it comes to memtary rest in 6m/s. The Diameter of wheel is?

OpenStudy (experimentx):

the find the angular speed by w = dx/dt put t=2, and get angular velocity ... use relation wr = s => w*r = 6 to find r.

OpenStudy (anonymous):

W = 4 - 4t at t=2 4-8 = -4 Wr = V r = V/W = 6/4 = 3/2 d = 3

OpenStudy (experimentx):

seems okay ... what's the answer?

OpenStudy (anonymous):

i dont knw but the options are 0.5 0.75 1.5 2